A vendor makes up a 20-kilo mixture of peanuts and green peas to come up with a mixture that cost Php 54 per kilo. If the peanut cost Php 60 per kilo and the green peas cost Php 45 per kilo, how many kilos of each kind did she use?
Answer by mananth(16946)

A vendor makes up a 20-kilo mixture of peanuts and green peas to come up with a mixture that cost Php 54 per kilo. If the peanut cost Php 60 per kilo and the green peas cost Php 45 per kilo, how many kilos of each kind did she use?
...
peanuts Php 60 --------x kilos
green peas Php 45------20-x kilos
Mixture ----Php=54-----20 kilos
...
60x+45(20-x)=54*20
60x+900-45x=1080
15x+900=1080
-900
15x=1080-900
15x=180
/15
x= 12 kilos ----------- peanuts.
Balance 8 kilos will be green p[eas.
...
m.ananth@hotmail.ca